OpenStudy (anonymous):

Evaluate: sqrt(0.16)=

OpenStudy (anonymous):

.4

jimthompson5910 (jim_thompson5910):

sqrt(0.16)=sqrt(16/100)=sqrt(16)/sqrt(100)=4/10=0.4 So sqrt(0.16)=0.4